A plane leaves Denver heading due north at 500 mph. Simultaneously, another
plane leaves Denver traveling due east at 1200 mph. After how many minutes
will the planes be 650 miles apart?

Answer by stanbon(57410)

A plane leaves Denver heading due north at 500 mph. Simultaneously, another
plane leaves Denver traveling due east at 1200 mph. After how many minutes
will the planes be 650 miles apart?
-------------------
Draw the picture on a coordinate system; use the y-axis for north
and the x-axis for east.
---------------------------
Let the time in hours be "x" for both planes
North distance = 500x miles
East distance = 1200x miles
-------------------------------
Use Pytagoras with a hypotenuse of 650 miles.
EQUATION:
(500x)^2 + (1200x)^2 = 650^2
250000x^2 + 1440000x^2 = 650^2
1690000x^2 = 650^2
x^2 = 0.25
x = 0.5 hours
------------------
Time = 30 minutes
========================
Cheers,
Stan H.